### v1

# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

Abstract
The safety and reliability of Automated Driving Systems (ADSs) must be validated prior to large-scale deployment. Among existing validation approaches, scenario-based testing has been regarded as a promising method to improve testing efficiency and reduce associated costs. Recently, the emergence of Large Language Models (LLMs) has introduced new opportunities to reinforce this approach. While an increasing number of studies have explored the use of LLMs in the field of automated driving, a dedicated review focusing on their application within scenario-based testing remains absent. This survey addresses this gap by systematically categorizing the roles played by LLMs across various phased of scenario-based testing, drawing from both academic research and industrial practice. In addition, key characteristics of LLMs and corresponding usage strategies are comprehensively summarized. The paper concludes by outlining five open challenges and potential research directions. To support ongoing research efforts, a continuously updated repository of recent advancements and relevant open-source tools is made available at: https://github.com/ftgTUGraz/LLM4ADSTest.
Index Terms:
Generative AI, simulation test, safety assessment, automated vehicle, literature review.I Introduction
Before the deployment of Automated Driving Systems (ADSs), their safety and reliability must be rigorously validated, a process traditionally requiring billions of miles on-road driving by Automated Vehicles (AVs) [1]. However, conventional mileage-based on-road testing has been deemed impractical due to its high costs and time-intensive nature. To address these limitations, the PEGASUS project [2] introduced the scenario-based testing approach, which enhances validation efficiency by subjecting ADSs to simulated and virtual driving environments. The effectiveness of this method has been substantiated, with simulation-based testing recongnized as a major contributor to the performance improvements of AVs by Waymo [3]. Consequently, the exploration of scenario-based testing has emerged as a crucial area of research.
Meanwhile, rapid advancements have been observed in the field of Artificial Intelligence (AI). Among the most prominent developments, Large Language Models (LLMs) have emerged as state-of-the-art AI systems capable of understanding and generating human language, and of performing a wide range of tasks without task-specific training [4, 5]. In recent years, researches at the intersection of LLMs and Automated Driving (AD) have gained increasing attention. As shown in Figure 1, subplot 1a illustrates the increasing volume of research publications related to LLMs and AD topics individually, while subplot 1b highlights a noticeable rise in studies specifically addressing the application of LLMs in the AD domain. This trend underscores a rapidly growing research direction focused on applying LLMs in development of ADSs.
Existing surveys have reviewed various aspects of LLMs, including their development process [6, 7, 8], reasoning capabilities [9] and prompting strategies [10, 11], as well as privacy-related concerns [12, 13, 14]. In parallel, several surveys have addressed scenario-based testing methodologies [15, 16, 17, 18, 19, 20, 21, 22, 23], and others have examined the application of LLMs in AD and related domains [24, 25, 26, 27, 28]. However, a systematic review specifically focusing on how LLMs are utilized within scenario-based testing of ADSs remains absent.
To the best of the authors’ knowledge, this work presents the first comprehensive survey that systematically examines how LLMs are integrated into various phases of scenario-based testing of ADSs, offering a broad perspective on emerging methodologies. The main contributions of this work are outlined as follows:
A taxonomy is established to classify existing approaches based on the testing phases and the specific roles assigned to LLMs.
The types of LLMs, their usage strategies, associated tasks, and the simulators integrated for scenario-based testing are systematically summarized.
Five key challenges are identified, and corresponding future research directions are proposed to advance the application of LLMs in this domain.
The remainder of this work is organized as follows: Section II reviews recent survey studies categorized by relevant topics. Section III introduces key terminologies used throughout the paper. Section IV presents a detailed analysis of specific tasks performed by LLMs. Section V summarizes the types of LLMs and their corresponding adoption strategies using tabular representations. Section VI briefly outlines the application of LLMs in industrial contexts. Section VII identifies five major challenges and discusses potential research directions. Finally, Section VIII presents the concluding remarks of this survey.
II Related Survey
This section provides an overview of recent relevant survey studies. It is organized into four categories: general surveys on LLMs, surveys focusing on scenario-based testing, surveys covering the use of LLMs in AD, and surveys addressing LLM applications in other domains.
II-A Survey of LLM
Recent surveys have examined the development and evolution of LLMs and their multimodal extensions. Naveed et al. [6] and Hadi et al. [7] offer comprehensive overviews of LLM architectures, training strategies, practical applications, and associated challenges. Zhang et al [8] systematically analyze 126 multimodal LLMs (MLLMs), with detailed comparison of model design and benchmark performances.
In addition to model development, several surveys have explored key functional aspects of LLMs. Su et al. [9] review efficient reasoning strategies, categorizing them into model-based, output-based, and prompt-based approaches. The domain of prompt engineering has been addressed by Sahoo et al. [10] and Chen et al. [11], who examine various methods, datasets, applications, and associated security implications.
Privacy preservation remains a critical limitation to broader deployment of LLMs. To address this concern, multiple surveys have investigated privacy-related challenges. Hu et al. [12] discuss the application of differential privacy in natural language models. Neel and Chang [13] provide a broader overview of privacy issues in LLMs, while Edemacu and Wu [14] focus specifically on privacy concerns related to in in-context learning and prompting mechanisms.
II-B Survey of Scenario-Based Testing
Scenario-based testing has been widely recognized as a critical methodology for assessing the safety and reliability of ADSs. A number of studies have reviewed this approach from various perspectives. Nalic et al. [15] and Zhang et al. [16] conduct extensive reviews encompassing over 80 publications and proposed taxonomies to classify scenario generation techniques. Zhong et al. [17] focus on scenario-based testing utilizing high-fidelity simulation platforms, whereas Ding et al. [18] categorize generation methods into data-driven, adversarial, and knowledge-based categories.
Beyond scenario generation, some studies have extended their scope to include safety assessment frameworks. Riedmaier et al. [19] advocate for the integration of formal verification methods within scenario-based testing. Similarly, Wishart et al. [20] review current verification and validation practices to support the development of SAE standards. Further, Alghodhaifi et al. [21] and Tang et al [22] investigate evaluation efficiency, emphasizing the role of accelerated and AI-based techniques as valuable complements to scenario-based approaches. Finally, Mihalj et al. [23] explore the interaction between ADSs and physical infrastructure, offering perspectives on how environmental factors can influence the design of test scenarios.
II-C Survey of LLM for Automated Driving
Recent surveys investigate the integration of LLMs and their multimodal extensions into ADSs, highlighting their potential to enhance perception, reasoning, and decision-making capabilities. Within a particular focus on Vision-Language Models (VLMs), Yang et al. [29] and Zhou et al. [30] review their applications across key functions such as perception, planning, control, and data generation, while underscoring their open-world understanding capabilities and the challenges involved. In a complementary study, Cui et al. [31] systematically review the development, tools, datasets, and challenges of applying MLLMs in AD and map systems, with the aim of informing future research and practical implementation.
Broader discussions are provided by Fourati et al. [32] and Li et al. [33], who examine the deployment of LLMs, VLMs and MLLMs within both modular and end-to-end system architecture, focusing on their structural designs, deployment strategies, and prospective research directions. At the system-level, Gan et al. [34] offer a comprehensive review of LLM applications in intelligent transportation systems, with an emphasis on scalable and efficient implementation.
II-D Survey of LLM for Miscellaneous Domains
Beyond the domain of AD, LLMs are extensively reviewed across a range of application areas. In the context of software testing, Braberman et al. [24] propose a taxonomy for LLM-based verification tasks, while Zhou et al. [25] outline frameworks and challenges associated with intelligent system testing. Yu et al. [26] survey alignment algorithms for MLLMs, and Luo et al. [27] review visual foundation models for road scene understanding. In the context of urban innovation, Xu et al. [28] examine the integration of generative AI models with urban digital twins, emphasizing their potential to support smart city management.
III Terminology
III-A Scenario
III-B Scenario Abstraction Level
According to [37, 38], scenarios are organized into four levels of abstraction: 1) Functional scenarios represent high-level traffic situations described in natural language using domain-specific terms. 2) Abstract scenarios are defined by transforming natural language descriptions into formats suitable for machine interpretation, often supported by ontologies. 3) Logical scenarios define the parameter ranges of the state values used for scenario representation. 4) Concrete scenarios use specific state values to ensure their reproducibility and to enable test methods to execute the scenario.
IV LLM applications by phase
for tree=
grow=east, parent anchor=east,
child anchor=west,
edge=ieeecyan, line width=1pt,
edge path=
[…]
rounded corners, draw, align=center, font=,
minimum size=3mm,
for children=
[Section IV LLM Applications by Phase, fill=ieeecyan!50, draw=none[IV-E ADS Assessment, fill=ieeecyan!50, name=ADSAssessment, draw=none
Safety performance evaluation [39, 40]
Intelligence level evaluation [41]
[IV-D Test Execution, fill=ieeecyan!50, name=TestExecution, draw=none
[Scenario optimization, fill=ieeecyan!50, draw=none
Syntax correction [42, 43, 44, 45]
Fidelity refinement [46]
[Simulation setup automation, fill=ieeecyan!50, draw=none
[39, 47, 48, 49]
[Anomaly detection, fill=ieeecyan!50, draw=none
[50]
[IV-C Scenario Selection, fill=ieeecyan!50, name=ScenarioSelection, draw=none
[Realism assessment, fill=ieeecyan!50, draw=none
[51, 52]
[IV-B Scenario Generation, fill=ieeecyan!50, name=ScenarioGeneration, draw=none
[LLM as an executable scenario generator, fill=ieeecyan!50, draw=none
Template filling [44]
End-to-end generation [43, 46, 53, 54]
Hybrid generation [42, 47, 48, 49, 55, 56, 57, 58]
[LLM as a standardized format generator, fill=ieeecyan!50, draw=none
OpenSCENARIO [55]
AVUnit [59]
[LLM as an intermediate format generator, fill=ieeecyan!50, draw=none
Driving policies [60, 61, 62]
Scenario elements [63, 64, 65, 66]
Functional scenario [67]
Abstract scenario [68]
Logical scenario [69]
[LLM as a data interpreter, fill=ieeecyan!50, draw=none
Accident report [70]
Domain-specific knowledge [71]
Naturalistic driving data [72, 73, 74, 75]
[LLM as a human-machine interface, fill=ieeecyan!50, draw=none
Structured scenario representation [76, 77, 39, 78, 79, 80]
Loss function [81]
Executable code [82, 83, 84]
[IV-A Scenario Source, fill=ieeecyan!50, name=ScenarioSource, draw=none
[Data retrieval, fill=ieeecyan!50, draw=none
Video [85]
Image [86, 87, 88]
[Data labeling, fill=ieeecyan!50, draw=none
[89]
[Data enrichment
through harzard analysis, fill=ieeecyan!50, draw=none
HARA [90, 91, 92]
STPA [93, 94, 95]
[Data enrichment, fill=ieeecyan!50, draw=none
Driving trajectory [96, 97, 98]
Future driving video [99]
Photo-realistic driving scene [100]
The scenario-based testing workflow, as illustrated in Figure 2, is initiated with scenario sources, which may include synthesized data, naturalistic driving measurements, or human expert knowledge. These sources are utilized to generate test scenarios, which are subsequently organized and stored within a scenario database. Scenario selection is then performed according to predefined criteria, followed by execution on simulation platforms such as CARLA [101] to evaluate the performance of ADSs. Finally, test results are analyzed and reported in the ADS assessment phase.
In this section, the application of LLMs in scenario-based testing within different testing phases are examined, along with corresponding implementation details. An overview of the research structure for this section is presented in Figure 3.
IV-A Scenario Source
Scenario-based testing of ADSs heavily relies on data, which serve as the basis for generating test scenarios. As a result, considerable attention has been directed toward scenario sources to support this process. As noted in [15, 19], these source can be classified into data-based or knowledge-based categories. Data may be obtained from real-world driving environments or synthetically generated using real-world contexts and expert input. Surveyed literature indicates that LLMs are primarily employed for tasks such as data enrichment, labeling, and retrieval. The implementation strategies associated with these applications are described below.
IV-A1 Data Enrichment
The acquisition of real-world driving data often falls short of meeting the evolving demands of ADSs, primarily due to the high costs and time-intensive nature of data collection and annotation. As a result, data synthesis has become an essential component for ADS testing [102]. LLMs have been employed to generate driving trajectories based on natural language instructions. In [96], LLMs are utilized for both trajectory-conditioned language generation and language-conditioned trajectory synthesis. In [97], right-hand driving trajectories are synthesized by LLMs using left-hand driving data as a reference. In [98], LLM is used to translate brief descriptions of vehicle interactions into realistic driving behaviors and trajectories by incorporating human-like driving logic.
Jia et al. [99] further expand data enrichment by combining an LLM with a video diffusion model to generate future driving scenes in video format. This enables trajectory generation and visual simulation based on natural language–driven vision–action pairs. Furthermore, LLMs have been adopted for ground truth generation; for example, Wei et al. [100] propose ChatSim, a framework that enables the creation of editable, photo-realistic 3D driving scenes through natural language commands. Within this framework, LLMs are used to facilitate user interaction and to enable the integration of external digital assets.
IV-A2 Data Enrichment Through Hazard Analysis
Hazard Analysis and Risk Assessment (HARA) and Systems Theoretic Process Analysis (STPA) are originally developed for functional safety and system-level hazard analysis, respectively. These methods are also commonly applied within the scenario-based approach described in ISO 21448 (cf. [103, 104, 16]) to systematically identify hazardous scenarios, gain insights into potential hazards (cf. [105, 106, 107]) and guide the generation of high-risk scenario seeds (cf. [108, 109, 110]).
HARA automation using LLMs has been explored to improve the efficiency and accuracy of safety analysis in ADSs. At Qualcomm [90], LLMs enable the identification of 20% more hazardous scenarios in an Autonomous Emergency Braking (AEB) case study compared to manual methods. Volvo Cars researchers investigate LLMs to generate scenarios and malfunctions from functional descriptions using a general framework, without relying on standardized databases [91], thereby demonstrating their potential for domain-independent safety analysis. A follow-up study [92] applies LLMs to generate safety requirements and detect inconsistencies, which enhances efficiency in safety engineering while highlighting the need for expert validation to ensure accuracy.
STPA automation using LLMs has also been investigated. Charalampidou et al. [93] used ChatGPT-4 to generate loss scenarios for an unmanned aerial vehicle rescue system, although the results require human validation. Similarly, Diemert et al. [94] introduce a cooperative framework where LLMs assist in identifying hazards, though expert oversight is required due to inaccuracies. Qi et al. [95] extended this to automotive and energy systems, assessing human-LLM collaboration schemes.
IV-A3 Data Labeling
Manually annotate data is both costly and labor-intensive, thereby necessitating automated solutions to enhance the scalability and efficiency of dataset development [102]. To address this challenge, LLMs have been utilized to partially substitute human effort in the annotation process. Chen et al. [89] introduce the first benchmark for evaluating Large Vision-Language Models (LVLM) in the context of self-driving corner cases. In their work, LLMs are employed to automatically generate annotations in JSON111https://www.json.org/json-en.html format, which were subsequently verified through manual inspection to ensure data accuracy and quality.
IV-A4 Data Retrieval
Effective scenario generation requires the retrieval of relevant information from large-scale, heterogeneous datasets, which often include multimodal data in varying formats and structures. This inherent complexity presents challenges for unified processing and has traditionally been managed through rule-based methods, such as Structured Query Language (SQL) queries (cf. [111]), which demands significant manual effort. Recent studies have instead employed LLMs to enable more intuitive and flexible data retrieval, thereby reducing the reliance on predefined rules and manual querying.
Video retrieval efficiency has been demonstrated to improve through the application of LLMs. In the study by Knapp et al. [85], MLLMs are integrated with a vector database to enable natural language querying of driving logs. Scenario descriptions were generated from sensor data and video content, allowing users to navigate large-scale datasets without relying on SQL-based queries.
Image retrieval efficiency has also been improved through the application of LLMs. In [86], three LVLMs are employed to perform image retrieval based on natural language queries. Their performance is evaluated using precision-based metrics on a benchmark dataset containing diverse driving scenarios with varying scene complexity, weather conditions, and traffic levels. Additionally, Rigoll et al. [87] propose an annotation-free, object-level image retrieval approach for AD datasets by combining panoptic segmentation with a VLM to support natural language queries. While this approach enhances data accessibility without manual labeling, it is limited to single-object recognition and performs poorly in complex, dynamic scenes involving multiple objects. To overcome these limitations, Tang et al. [88] incorporate LLMs and Bird’s Eye View (BEV) representations to enhance text-to-scene retrieval. This approach is evaluated on nuScenes-Retrieval dataset, an extension of nuScenes dataset [112] with enriched and diverse textual annotations. Experimental results show that it achieves state-of-the-art performance, clearly outperforming existing baseline methods.
IV-B Scenario Generation
This section reviews previous studies that have employed LLMs for scenario generation. To enable execution within simulation platforms such as CARLA [101], the generated scenarios must be converted into standardized formats, such as ASAM OpenSCENARIO (XOSC) [113]. Given the inherent complexity of scenario generation, the process is typically decomposed into several sub-steps. When LLMs are utilized, their contributions are generally restricted to specific stages of the workflow rather than spanning the entire process. An overview of the functional roles played by LLMs throughout the scenario generation pipelines is illustrated in Figure 4, and the subsequent discussion is organized accordingly.
IV-B1 LLM as a Human-Machine Interface
In this subsection, LLMs are employed to translate natural language inputs provided by the user into structured information that facilitates downstream scenario generation, as illustrated in step ① of Figure 4. Specifically, LLMs have been used to interpret user input into structured scenario representations [76, 77, 39, 78, 79, 80], loss function formulations [81], and executable code [82, 83, 84]. The primary strategies adopted in these studies are schematically summarized in Figure 5.
Structured scenario representations derived from user inputs have been employed in various studies to support scenario generation. In [76], such structured information is employed to create executable files through a diffusion-based model. Similarly, in [77] and [39], it is used to trigger Python scripts for automated file generation. Building upon [77], Zhou et al. [114] evaluate the performance of six different LLMs in interpreting motorway functional scenarios, thereby offering a benchmark for model selection. In [78], structured representations are used as inputs to a transformer model for generating vehicle trajectories. In [79], LLMs convert natural language inputs into scenario elements, which are subsequently transformed into simulation-executable files. A unified diffusion-based framework is introduced in [80], where LLMs generate Proto222https://protobuf.dev/ constraints from user input to enable language-driven control of scene initialization and closed-loop simulation.
Loss functions and executable code have also been generated from natural language inputs using LLMs. In [81], a loss function is derived from user-provided descriptions and integrated into a diffusion model to facilitate executable file generation. Code generation from natural language inputs is further demonstrated in [82, 83]. In [82], the generated code, which includes interaction, vehicle, and map modules, is utilized by a transformer to synthesize vehicle trajectories. In [83], the generated code is applied to call functional libraries for producing vehicle trajectory data. In a related study, Nguyen et al. [84] employ a LLM to parse user language into driving behavior code, which is then used to guide reinforcement learning for generating diverse synthetic driving scenarios.
IV-B2 LLM as a Data Interpreter
This subsection reviews studies in which LLMs have been employed to interpret various data sources into natural language representations. This process is illustrated in step ② of Figure 4. Scenario-relevant information has been extracted from diverse inputs, including accident reports [70], domain-specific knowledge [71], and naturalistic driving data [72, 73, 74].
Accident reports have been interpreted by LLMs to support scenario reconstruction. In the study by Guo et al. [70], a framework is proposed in which textual accident reports are parsed into a structured JSON format, which is subsequently processed by a constraint solver to generate waypoints.
Additionally, domain-specific knowledge has been extracted using LLMs to aid scenario generation. Tang et al. [71] introduce a method in which technical documents are interpreted and transformed into structured ontological elements, thereby enabling automated ontology construction.
Naturalistic driving data have also been utilized to support sceario generation through LLM-based interpretation. Mei et al. [72] employ LLMs to identify the most threatening attacker in each scene from the Waymo Open Dataset [115], enabling the creation of high-risk scenarios for improving the testing and training of ADSs. In [73], LLMs are applied to interpret BEV maps derived from the nuPlan dataset [116], with the extracted spatial information subsequently used to guide a diffusion model for trajectory generation. Tian et al. [74] integrate reinforcement learning and LLMs to optimize safety-critical driving scenarios within the HighwayEnv [117] simulation environment. Extending these efforts into real-time application, Mei et al. [75] focus on online scenario generation by using retrieval-augmented LLMs to infer dangerous driving behaviors and synthesize adversarial trajectories based on historical states.
IV-B3 LLM as an Intermediate Format Generator
The direct generation of simulator-executable scenario files is inherently complex. To manage this complexity, the process is typically divided into multiple stages, with the outputs of each referred to in this study as intermediate formats. These formats acts as transitional representations that can be subsequently transformed into executable scenario files, as illustrated in step ③ of Figure 4. Unlike steps ① and ②, which primarily rely on either data or knowledge sources, step ③ integrates both.
A common application of LLMs in this context involves the generation of driving policies based on natural language inputs. In [60] and [61], user prompts are combined with environmental context and processed to derive driving strategies, which are subsequently used to inform either a driver model [60] or an auto-regressive trajectory generation model [61]. Similarly, Wei et al. [62] demonstrate the use of LLMs to interpret linguistic descriptions for trajectory planning, thereby guiding behavior of autonomous agents.
In addition, LLMs have been employed to synthesize various scenario elements. In [63], an approach is presented in which maps and vehicle assets are generated from textual descriptions. Waypoints are selected using a VLM, followed by trajectory generation using a diffusion model. In [64], critical corner cases are produced by integrating user language inputs, failure records, and scenario databases content. This methodology is further extended in [65], where scenario mutation, prior test feedback, and expert knowledge are incorporated to enhance scenario diversity. Similarly, in [66], descriptive inputs and datasets are processed to generate scenario configurations and associated parameter sets.
LLMs have also been employed to derive functional [67], abstract [68], and logical scenario [69] representations. In [67], a MLLM is applied to analyze accident videos, generate narrative descriptions, and identify relevant objects, thereby enabling the transformation of raw video content into functional scenario representations. In [68], traffic regulations are parsed to produce abstract scenarios with defined syntax, which are subsequently converted into executable scripts through code generation. Furthermore, in [69], accident reports are interpreted to extract logical scenarios, which are then instantiated as concrete scenarios using a search-based algorithm.
IV-B4 LLM as a Standardized Format Generator
This subsection reviews studies in which LLMs are employed to generate standardized scenario formats from intermediate representations, rather than directly from scenario source, as illustrated in step ④ of Figure 4. In the study by Zorin et al. [55], accident data related to ADSs are collected from online sources and converted into key textual descriptions using Python scripts. LLMs are then utilized in conjunction with predefined templates to produce scenario files compliant with the XOSC standard. Similarly, Tian et al. [59] employ Large Multimodal Models (LMMs) to generate safety-critical scenarios from non-accident traffic videos. Optical flow data and Chain-of-Thought (CoT) reasoning are used to construct abstract representations, which are subsequently transformed into executable programs. The generated scenarios are validated through a dual-layer optimization framework, and the resulting trajectories are encoded using AVUnit [118].
IV-B5 LLM as an Executable Scenario Generator
This subsection reviews studies in which LLMs are utilized to generate executable scenarios directly from scenario sources, as illustrated in step ⑤ of Figure 4. The literature indicates that LLMs have primarily been applied to produce scenarios in standardized formats, including XOSC [113] (cf. [53, 55, 119]), SCENIC [120] (cf. [54, 46, 43, 56, 49]), SUMO XML333https://sumo.dlr.de/docs/Networks/PlainXML.html (cf. [42, 44, 48]), and AVUnit [118] (cf. [57]), as well as in user-defined formats (cf. [47, 58]).
A range of techniques has been adopted in the reviewed studies to generate simulator-executable files. These approaches can be categorized into three types: template filling, end-to-end generation, and hybrid generation, depending on the degree of reliance on external tools. This classification is illustrated in Figure 6.
Template Filling
Template filling refers to the process in which LLMs populate predefined scenario template with specific parameters in the to generate executable files. Within this category, the study presented in [44] is representative. In this work, Güzay et al. propose a method for generating SUMO scenario based on predefined templates. Structured prompts are designed to guide GPT-4 in producing XML files according to user-defined variables, such as the number of intersections and the length of road segments.
End-to-End Generation
End-to-end generation refers to the direct mapping of natural language or multimodal inputs into executable simulation files, without relying on predefined templates. This paradigm is first explored by Wang et al. [53], who prompt general-purpose LLMs to generate XOSC files. To enhance controllability, Miceli-Barone et al. [43] propose a dialogue-based system in which SCENIC programs are incrementally constructed through iterative user interactions. Extending this approach to the visual domain, Miao et al. [46] introduce ScriptGPT, which translates dashcam videos into SCENIC scripts using prompt-engineered video-language models. To address the issue of structural correctness, Elmaaroufi et al. [54] combine compositional prompting with compiler feedback to generate probabilistic scenarios from crash reports.
Hybrid Generation
Hybrid generation approach combines template-filling and end-to-end generation, enabling greater flexibility while maintaining scenario standardization. In [55], Zorin et al. integrate real-world traffic incident data scraped from websites with LLM-based scenario completion methods to bridge open-source intelligence and executable simulation file generation. Yang et al. [49] introduce SimCopilot, which translates natural language descriptions of object interactions into simulation-ready code using a language-to-interaction dataset, thereby balancing high-level abstraction with detailed execution. Lu et al. propose OmniTester [42] and AutoScenario [48], both of which incorporate prompt engineering, Retrieval-Augmented Generation (RAG), and external simulation tools (e.g., SUMO, CARLA) to enhance the realism, diversity, and controllability of generated scenarios. Similarly, Zhang et al. [56] propose ChatScene, which decomposes textual inputs into subcomponents and retrieves relevant domain-specific code snippets to generate safety-critical scenarios. Aiersilan et al. [47] present AutoSceneGen that leverages in-context learning to transform high-level textual descriptions into simulator-compatible configuration files, enabling batch generation of safety-critical scenarios in CARLA.
IV-C Scenario Selection
Scenario selection is performed to efficiently filter concrete scenarios from a database, with the aim of identifying cases that may reveal potential issues in ADSs [19]. As a recently explored approach, realism assessment emphasizes evaluating the degree to which generated scenarios align with real-world conditions, thereby allowing only realistic cases to be selected in order to improve test validity. LLMs have been applied to support this evaluation process.
Two representative studies have investigated the use of realism assessment for scenario selection. In the study by Wu et al. [51], LLMs are employed to evaluate the consistency of driving trajectories with real-world conditions. This is accomplished by applying standardized prompts to assess perturbed variants of the DeepScenario dataset [121], resulting in a robustness score that quantifies scenario realism. In a related effort, Fu et al. [52] combine a diffusion model with a VLM to generate, narrate, and interpret realistic driving videos. The realism of the generated outputs is verified to enhance scene understanding. Experiments conducted on the Waymo Open Dataset [122] demonstrate the effectiveness of this approach in advancing VLM applications for AD.
IV-D Test Execution
In scenario-based testing, concrete scenarios are executed within test environments such as real-world roads, close-track, or simulation platforms, each supporting varying levels of X-in-the-loop (XiL) integration [123, 19]. The studies reviewed in this work primarily focus on the use of LLMs in fully simulated environment. In these contexts, LLMs are employed to dynamically adjust testing parameters, thereby enabling self-adaptive testing procedures.
IV-D1 Anomaly Detection
ADSs are susceptible to failures resulting from system-level deficiencies, which are commonly referred to as anomalies. Detecting such anomalies requires advanced reasoning capabilities. In this context, LLMs have been utilized to identify perception system anomalies in real time during test execution phase. Elhafsi et al. [50] apply OWL-ViT [124], a vision transformer model designed for object-level visual understanding, to extract visual feature, and GPT-3.5 for reasoning to detect anomalies in ADSs within the CARLA simulation environment. Their approach outperforms traditional Out-Of-Distribution (OOD) detection methods while offering interpretable analysis, although limitations remain due to the reliability of object detection and LLM inference.
IV-D2 Simulation Setup Automation
The configuration of simulation environments remains a complex and expertise-intensive task. Recent advancements in LLMs offer a promising solution for automating simulation setup by translating natural language inputs into executable configurations, a direction explored by several recent studies. In [39, 47] and [48], LLMs are employed to modify traffic environments, generate traffic scenarios, and configure simulation files for platforms SUMO [39] and CARLA [47, 48]. These approaches support real-time adjustments, enhance scenario diversity, and enable end-to-end automation without manual intervention. Building upon this research direction, Yang et al. [49] construct virtual traffic scenes in the LGSVL simulator based on natural language descriptions and introduce a language-to-interaction dataset introduced to support benchmarking and further development in simulation setup automation.
IV-D3 Scenario Optimization
During the execution of concrete scenarios in physical simulators, error messages may occur due to syntax errors in scenario definitions. Traditionally, the correction of such errors has been performed manually and offline. To improve efficiency, Lu et al. [42] utilize a LLM integrated with a RAG module to iteratively validate and refine scenarios based on simulation feedback from CARLA and SUMO. Similarly, in [43], an LLM is used to automatically correct and update SCENIC code in response to both simulation results and user feedback, until successful execution is achieved in CARLA. In [44], GPT-4 is applied to parse error messages generated by SUMO and to modify the simulation files accordingly. Beyond syntax correction, LLMs have also been utilized for generating and refining closed-loop control code. In [45], an LLM is leveraged to generate and iteratively refine control code from natural language descriptions based on simulation feedback in Esmini [125], thereby forming a closed-loop workflow integrating code generation, simulation, and correction.
Furthermore, if simulated scenarios are not consistent with real-world driving data, the resulting simulations may lack reliability. To refine scenario fidelity during execution, Miao et al. [46] utilize LLM to perform similarity analysis between simulated scenarios and dashcam crash video. In case of mismatch, the LLM iteratively refines the SCENIC script by incorporating feedback from a similarity metric. Scenario elements are updated until the simulated scenario closely approximates the original crash event.
IV-E ADS Assessment
In this phase, the performance of ADSs is quantified using key metrics, with Time-To-Collision (TTC) [126] being one widely adopted example. These statistical indicators are typically included in simulation test report. To automate this process, LLMs have been employed to generate simulation reports. In [39], a LLM is used to interpret SUMO output files, extract relevant metrics such as traffic density, travel time, and emissions, and synthesizes analytical summaries to support efficient evaluation and comparison of simulation outcomes.
Beyond simulation, LLMs have also been applied to the analysis of real-world ADS accidents. Xu et al. [40] explore the potential of LLMs to generate legal explanations for ADS-related accident cases. The limitations of the models are evaluated, and improvements are proposed through domain-specific adaption and enhanced contextual reasoning.
In addition to safety assessment, LLMs have been explored as tools to evaluate the intelligence level of ADSs. You et al. [41] employ LLMs to perform hierarchical assessments by using CoT prompting combined with RAG. The models simulate human-like reasoning across multiple decision-making levels, and their outputs are validated through CARLA simulations and human evaluators.
V Model selection and adaption
LLMs have begun to play a crucial role in scenario-based testing of ADSs, with numerous LLMs proposed from different organizations. Table I and Table II provide a comprehensive summary of the LLMs used in the reviewed studies, categorized according to various attributes.
Country | Organization | Model | Open- source | Testing Phase | Usage Strategy | Task | Related Work | Integrated Simulator |
USA | OpenAI | GPT-4 | - | ① | ❺ | STPA | [93] | - |
❹ | HARA | [91] | - | |||||
❸❹ | Visual scene editing | [100] | - | |||||
❷❹ | HARA | [92] | - | |||||
❹ | HARA | [95] | - | |||||
②⑤ | ❷❸❹ | Functional scenario generation | [47] | CARLA | ||||
❷❸❼ | Scenario generation | [42] | SUMO | |||||
❷❸❹❼ | DSL scenario generation | [79] | - | |||||
❸ | Scenario generation | [64] | - | |||||
AVUnit generation | [57] | LGSVL | ||||||
❹ | Accident report interpretation | [70] | LGSVL | |||||
② | AVUnit generation | [59] | LGSVL | |||||
Driving behavior generation | [77] | Esmini | ||||||
❷❹❼ | Functional scenario interpretation SCENIC generation | [56] | CARLA | |||||
❶❹ | XOSC generation | [55] | - | |||||
Cost function generation | [76] | - | ||||||
❶ | Scenario encoder generation | [82] | - | |||||
❷❸ | Multimodal data interpretation | [48] | CARLA | |||||
Scenario generation | SUMO | |||||||
❶❹❼ | Driving behavior generation | [84] | - | |||||
❶❸ | Functional scenario interpretation | [78] | - | |||||
Scenario generation | [66] | - | ||||||
❷❹ | Loss function generation | [81] | - | |||||
DSL scenario generation | [69] | LGSVL | ||||||
SCENIC generation | [43] | CARLA | ||||||
Traffic rule interpretation JSON generation | [68] | CARLA | ||||||
SUMO configuration generation | [44] | SUMO | ||||||
❷❸❹ | User language interpretation | [62] | Isaac Gym Blender | |||||
Driving trajectory generation | ||||||||
⑤ | ❶❹ | Code generation | [45] | Esmini | ||||
GPT-4o | - | ②⑤ | ❷❸❹ | Functional scenario generation | [47] | CARLA | ||
② | ❷ | SCENIC generation | [46] | CARLA | ||||
❷❸❹ | JSON generation | [58] | CARLA | |||||
❶❹ | Accident video classification | [67] | - | |||||
GPT-4o-mini | - | ② | ❸❹ | Scenario generation | [65] | CARLA | ||
❶❹ | Accident video classification | [67] | - | |||||
GPT-4v | - | ① | ❹❻ | Visual grounding | [89] | - | ||
Text-davinci-003 | - | ⑤ | ❶ | Semantic anomaly detection | [50] | CARLA | ||
CLIP [127] | ① | ❹ | Image retrieval | [87] | - | |||
❶❹ | Scenario retrieval | [86] | - | |||||
GPT-3 | - | ① | ❹ | STPA | [94] | - | ||
④ | ❹ | Scenario realism assessment | [51] | |||||
GPT-3.5 | - | ② | ❶❷❸❹❼ | SCENIC generation | [54] | CARLA | ||
❶❷ | Domain knowledge distillation | [71] | - | |||||
❻ | XOSC generation | [53] | - | |||||
❺ | Trajectory generation | [83] | - | |||||
① | ❷❸ | Trajectory generation | [97] | - | ||||
Text-Embeeding -Ada-002 | - | ① | ❸❹ | Visual scene editing | [100] | - | ||
Anthropic | Claude-instant-1.2 | - | ② | ❶❷❸❹❼ | SCENIC generation | [54] | CARLA | |
LMSYS | Vicuna-7B-1.5 [128] | ① | ❶❹ | Driving video generation | [99] | - |
Country | Organization | Model | Open- source | Testing Phase | Usage Strategy | Task | Related Work | Integrated Simulator |
UW Madison | LLaVA-v1.5-7b [129] | ① | ❷ | Driving log retrieval | [85] | - | ||
LLaVA-v1.5-13b [129] | ② | ❹❻ | BEV interpretation | [73] | - | |||
BakLLaVA [130] | ① | ❶❹ | Scenario retrieval | [86] | - | |||
ByteDance | LLaVA-NeXT- Interleave [131] | ② | ❶❹ | Accident video classification Visual grounding | [67] | - | ||
LLaMA [132] | ②⑤ | ❻ | SCENIC generation | [49] | LGSVL | |||
LLaMA-2 [133] | ① | ❻❽ | Scenario retrieval | [88] | LGSVL | |||
④ | ❹ | Scenario realism assessment | [51] | - | ||||
LLaMA-2-7B [133] | ② | ❻ | XOSC generation | [53] | - | |||
USA | Meta | LLaMA 3.1 | ②⑤⑥ | ❹❼ | Road generation | [39] | SUMO | |
② | ❸❹ | Functional scenario generation | [72] | MetaDrive | ||||
LLaMA-3 [134] | ② | ❶❹ | XOSC generation | [55] | - | |||
LLaMA 3-8B | ② | ❹❻ | Driving behavior generation | [61] | - | |||
TinyLlama [135] | ① | ❻ | Trajectory interpretation | [96] | - | |||
CodeLlama [136] | ⑤ | ❶❹ | Code generation | [45] | Esmini | |||
Salesforce | BLIP-2 [137] | ① | ❶❹ | Scenario retrieval | [86] | - | ||
Sentence-T5 [138] | ② | ❷❹❼ | SCENIC generation | [56] | CARLA | |||
CodeGemma [139] | ② | ❷ | SCENIC generation | [45] | Esmini | |||
Gemma-7B [140] | ① | ❷ | Driving log retrieval | [85] | - | |||
Gemini 1.0 Ultra | - | ② | ❷❹❼ | DSL scenario generation | [80] | - | ||
China | Alibaba | Qwen-turbo | - | ② | ❶ | Driving behavior generation | [60] | CARLA SUMO |
Baidu | Yiyan | - | ② | ❻ | XOSC generation | [53] | - | |
Tsinghua Uni. | GLM-4-Air [141] | ② | ❸❹ | Scenario generation | [65] | CARLA | ||
DeepSeek | DeepSeek-Coder [142] | ⑤ | ❶❹ | Code generation | [45] | Esmini | ||
DeepSeek-R1 [143] | ⑤ | ❶❹ | Code generation | [45] | Esmini | |||
② | ❷❼ | Driving behavior analysis | [75] | - | ||||
DeepSeek-V3 | ② | ❷❼ | Driving behavior analysis | [75] | - | |||
France | Mistral AI | Mistral-7B-Instruct | ② | ❷❹ | Scenario generation | [74] | - | |
Mistral-7B | ④ | ❹ | Scenario realism assessment | [51] | - | |||
Mistral-7B | ④ | ❶❹ | Code generation | [45] | Esmini | |||
NA | NA | NA | NA | ① | ❺ | STPA | [90] | - |
NA | NA | ② | ❹ | XOSC generation | [119] | Esmini |
V-A LLM Origin
Figure 7 illustrates the distribution of LLMs applied in scenario-based testing of ADSs, categorized by organization and region. The majority of applied models originate from organizations based in the USA, with OpenAI leading by a substantial margin. In fact, the number of studies employing OpenAI models exceeds the combined total from all other organizations. China ranks second, with five models developed by three organizations, including recent contributions from DeepSeek. Within Europe, France is the only contributing country, represented by models primarily released by Mistral AI. These findings are further substantiated by the “Country” and “Organization” columns in Table I, and Table II.
Among the models adopted in the reviewed studies, the GPT series developed by OpenAI is the most frequently used, despite being closed-source. This is followed by the LLaMA series and the LLaVA series, both of which are open-source.
V-B LLM Usage Strategy
Although LLM usage strategies are generally applicable across domains, this section focuses on those specifically adopted in scenario-based testing of ADSs, aligning them with distinct testing phases. Rather than proposing a universal taxonomy, the aim is to categorize existing strategies observed in the reviewed literature. Definitions and classifications are presented in Table III.
Category | Subcategory | Definition |
Prompt Engineering | Zero-shot Learning | LLMs can perform complex reasoning tasks in a zero-shot manner without any task-specific examples [144]. |
Few-shot Learning | It refers to the models’ ability to perform tasks by conditioning on a few examples provided in the input prompt, without any fine-tuning [5]. | |
Chain-of-Thought | It is defined as a series of intermediate natural language reasoning steps that lead to the final output [145]. | |
Format Constraint | It refers to that LLMs output contents following predefined formats and relevant investigations have been introduced in. | |
Parameter Tuning | Full Fine-tuning | It refers to the process of updating all parameters of a pre-trained LLM to adapt it comprehensively to a new domain [146]. |
Low-Rank Adaption (LoRA) | Instead of updating all parameters, it adds small trainable layers while keeping the original model frozen, thereby reducing training cost and memory usage while maintaining strong performance on new tasks [147]. | |
Knowledge- Augmented Generation | Retrieval-Augmented Generation (RAG) | It integrates a retriever to fetch relevant documents from an external corpus with a generator that conditions on both the retrieved content and the original input to produce more informed and accurate responses [148]. |
Knowledge Graph Integration | Knowledge graphs are integrated to enhance LLMs’ reasoning and factual accuracy [149]. |
Figure 8 presents a heatmap illustrating the distribution of LLM usage strategies (❶–❽) across scenario-based testing phases (①–⑥). The analysis indicates that early testing phases, particularly scenario generation (②), exhibit the highest concentration of LLM applications. Among the various strategies, prompt engineering is the most frequently employed, with format constraints (❹) and few-shot learning (❷) being especially prevalent. Among these, format constraint stands out as the most commonly adopted strategy overall. In contrast, parameter tuning (❺ and ❻) and knowledge-augmented generation techniques (❼ and ❽) appear less frequently and are mainly confined to scenario generation (②). Minimal engagement is observed in ADSs assessment (⑥), suggesting that the integration of LLMs into the later-stage validation process remains limited and presents opportunities for future exploration.
V-C Tasks Performed by LLM
In Section IV-B, reviewed studies are classified based on the roles LLMs play in various tasks. To offer a more detailed understanding, specific tasks undertaken by LLMs are outlined in Table I and Table II.
V-D Integrated Simulator
Figure 9 illustrates the frequency of simulator usage in the reviewed studies. CARLA [101] is the most widely adopted platform, appearing in 15 studies. LGSVL [150] and Esmini [125] are each used in 7 studies, followed by SUMO [151], which is used in 5. Other simulators, including MetaDrive [152], CarMaker [153], and Isaac Gym [154], are each employed in only a single study. These results suggest a strong preference for open-source environment such as CARLA and LGSVL in the current research landscape, while the adoption of commercial remains limited.
VI Industrial perspective
Table IV presents an overview of how LLMs are adopted by industry in the context of scenario-based testing for ADSs. The organizations are categorized by region, with corresponding tasks and testing phases indicated.
In China, 51WORLD and Huawei adopt LLMs during the scenario generation phase for XOSC creation. In addition, Huawei applies LLMs during the scenario source phase for data labeling and retrieval. In Europe, Denotic, Luxoft, and IAV utilize LLMs in the scenario generation phase for XOSC generation, while dSPACE and Fraunhofer incorporate them into test automation workflows. Wayve employs LLMs in the scenario source phase for large-scale data generation. In the USA, Applied Intuition uses LLMs in the scenario generation phase for scenario construction.
Although numerous academic studies have explored the application of LLMs in scenario-based testing of ADSs, industrial adoption remains relatively limited. As shown by the relatively small number of documented cases in Table IV, current applications in industry are still at an early exploratory stage. Furthermore, existing use cases are primarily confined to a narrow set of tasks, such as XOSC generation and test automation, highlighting a noticeable gap between academic advancements and their translation into practical industrial deployment.
Region | Organization | Testing Phase | Task | Work |
China | 51WORLD | ② | OpenX Generation | [155] |
Huawei | ①② | Data Labeling Data Retrieval Scenario Generation | [156] | |
EU | Denotic | ② | OpenX Generation | [157] |
IAV | ② | OpenX Generation | [158] | |
Luxoft | ② | OpenX Generation | [159] | |
dSPACE | ⑤ | Test Automation | [160] | |
Fraunhofer | ⑤ | Test Automation | [161] | |
Wayve | ① | Data Generation | [162] | |
USA | Applied Intuition | ② | Scenario Generation | [163] |
VII Open challenges
The preceding sections have reviewed LLM applications across various phases of scenario-based ADSs testing. Despite notable progress, practical implementation remains challenging due to various constraints. This section outlines five key challenges, each discussed in the following sections.
VII-A LLM Hallucination and Output Variability
A key challenge in applying LLMs to scenario-based testing is their inherent response uncertainty, primarily in the format of hallucination and output variability. Hallucination represents to the generation of outputs that are either unfaithful to the input or factually incorrect [164, 165]. This issue arises across all testing phases, including physically implausible trajectories in data enrichment [97], annotation of non-existent in data labeling [89], invalid or non-executable code in scenario generation [56], and factually incorrect judgments in safety- or legality-critical assessments due to insufficient domain knowledge [40].
Output variability refers to nondeterministic behavior of LLMs, where identical prompts may yield divergent results. This unpredictability poses particular challenges in scenario generation, where minor variations can lead to cumulative structural errors [71]. For example, GPT-4 has been observed to generate inconsistent simulation configurations from identical prompts [44].
Several strategies have been proposed to mitigate these challenges. Huang et al. [165] categorize countermeasures into three levels: data, training, and inference and identify RAG as a promising approach. Elmaaroufi et al. [54] further propose using a secondary LLM as a verification agent to assess the output consistency with the given task description. Nonetheless, robust consistency guarantees remain an open problem.
VII-B Simulation Platform Integration
Another challenge in applying LLMs to scenario-based ADSs testing lies in their limited integration with simulation platforms, stemming from fragmented scenario formats, inconsistent interfaces, and constrained automation.
Fragmented scenario formats with unique syntax are adopted in the reviewed studies. Such heterogeneity often requires reworking prompts, templates, or conversion scripts. For example, studies [44, 53, 42] demonstrate how the same LLM (GPT-4) must be re-prompted to accommodate format differences.
Inconsistent interfaces, such as different Application Programming Interfaces (APIs) and import mechanisms across simulation platforms such as CARLA, SUMO, Esmini, and LGSVL, further hinder the integration. LLM-generated outputs often require translation or manual debugging, which reduces portability and limits automation.
Constrained automation has been demonstrated, although LLMs have shown promise in automating simulation setup. Core simulation elements like sensor configurations, vehicle dynamics, and environmental settings are still manually defined. Furthermore, real-time interaction between simulator and LLMs remains unrealized.
Developing a standardized intermediate representation could decouple LLM outputs from simulator-specific formats. Additionally, establishing feedback loops between simulation platforms and LLMs may further enable runtime iterative refinement.
VII-C Lack of Domain-Specific Models
General-purpose LLMs face notable limitations in scenario-based ADSs testing due to insufficient domain adaptation. This issue is exacerbated by the predominant reliance on prompt engineering over domain-specific fine-tuning, as well as the absence of authoritative ADSs datasets. For example, LLaMA generates invalid OpenSCENARIO syntax [53], while GPT-4 fails to interpret updated SUMO rules [44], revealing gaps in both semantic precision and temporal awareness.
These limitations propagate across testing pipelines: biased scenario selection [51], syntax validation failure (e.g., ISO 21448 misinterpretation [36, 45]), and unclear performance evaluation in safety-critical contexts [41]. A hybrid approach combining domain fine-tuning and RAG could mitigate these issues by linking LLMs to structured knowledge.
VII-D Insufficient Attention to Scenario Database
As discussed in previous sections, no studies have employed LLMs to construct scenario databases, which lags significantly behind other testing phases. This indicates a lack of systematic mechanisms for storing, managing, and reusing large volumes of generated scenarios. The construction of scenario databases should prioritize the development of standardized interfaces that integrate diverse data sources and convert them into machine-readable formats [19].
LLMs could assist in verifying the syntax and physical consistency of newly generated scenarios along with similarities, and performing de-duplication and quality assurance. Furthermore, LLM may support scenario analysis by comparing with regulations to generate coverage report.
VII-E Industrial Application Gap
While LLMs show promise for scenario-based testing, their industrial adoption remains limited due to several constraints: 1) Data privacy and security: Utilizing LLMs often requires transmitting sensitive data (e.g., vehicle trajectories) to third-party cloud services (e.g., OpenAI Enterprise444https://openai.com/chatgpt/enterprise/), which risks violating local data protection regulations. Although enterprise APIs provide contractual safeguards, they typically prohibit model fine-tuning, thus restricting task-specific performance. Alternatively, deploying open-source LLMs locally can preserve data privacy, but requires substantial infrastructure investment. 2) Explainability: According to ISO 21448 [36], scenario generation must be traceable. However, LLMs often lack transparency in their generation process and cannot justify why specific scenarios are produced, thereby creating gaps in safety-critical applications.
VIII Conclusion
This survey presents a structured overview of the application of LLMs in scenario-based testing of ADSs. By introducing a phase-based taxonomy, existing studies are systematically categorized according to roles LLM played in performed tasks in the testing pipeline. In addition, LLM types, usage strategies, and integrated simulators are quantitatively summarized, five key challenges that hinder LLM-enabled ADSs testing are identified.
The findings indicate that current applications are concentrated in early testing phases, particularly scenario generation, while downstream stages especially scenario database construction remain underexplored. Future researches should prioritize the development of standardized intermediate formats, hybrid approaches combining domain fine-tuning with knowledge-augmented generation, and closed-loop co-simulation frameworks that allow real-time interaction between LLMs and simulators.